OpenStudy (anonymous):

Give the general solution of the ODE: y'y = sqrt(1+y^2)cos(x)

OpenStudy (anonymous):

\[y \frac{ dy }{ dx } = \sqrt{(1+y^2} )\cos(x)\] \[\int\limits \frac{ y }{ \sqrt{(1+y^2} ) }dy = \int\limits\cos(x)dx\] \[\int\limits \frac{ 1 }{ \sqrt{(1+y^2} ) }d(1+y^2) = \int\limits \cos(x)dx\]

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2 } \sqrt{1+y^2} = \sin(x) + C)\]

OpenStudy (anonymous):

would that then become\[y = 2\sin(x) - 1 +c\] ?? @abtster

OpenStudy (anonymous):

constants and the like just change the constant

OpenStudy (anonymous):

c + 1 = c' etc

OpenStudy (anonymous):

ok that makes sense, thanks! @abtster

OpenStudy (perl):

nice solution

OpenStudy (perl):

@abtster I dont understand your integral. how did you get dy = d(y^2+1) , i dont understand that step

OpenStudy (anonymous):

@perl , \[ydy = \frac{1}{2}dy^2=\frac{1}{2}d()y^2+1)\] O ohh, my bad I forgot the fraction and some more flaws

OpenStudy (anonymous):

\[−\frac{1}{4}\sqrt{(1+y^2)}=\sin(x)+C\]

OpenStudy (anonymous):

\[\sqrt{1+y^2}=4sin(x)+C'\] \[1+y^2=(4sin(x)+C)^2\] \[y^2=(4sin(x)+C)^2-1\] \[y=\sqrt{(4sin(x)+C')^2-1}\] or \[y=-\sqrt{(4sin(x)+C')^2-1}\] there might be some trigonometry identities to simplify this, but I can't find it